# BPDec: Unveiling the Potential of Masked Language Modeling Decoder in BERT Model Pretraining

### v3

@@ 4.2 Performance on GLUE Tasks
  • AX (GLUE Diagnostic Task): Tests ability to understand linguistic phenomena such as logic, predicate-argument structure, and lexical semantics. The evaluation metric is typically overall accuracy across all diagnostic categories.
  • CoLA (Corpus of Linguistic Acceptability): Measures grammatical correctness. The metric used is Matthew’s correlation.
  • MNLI-m/mm (Multi-Genre NLI - Matched/Mismatched): Assesses the ability to predict textual entailment across different genres. Accuracy is the metric for both the matched (MNLI-m) and mismatched (MNLI-mm) sections.
  • MRPC (Microsoft Research Paraphrase Corpus): Focuses on identifying whether two sentences are paraphrases of each other. Accuracy is used as the main evaluation metric.
  • QNLI (Question NLI): Involves determining whether a context sentence contains the answer to a question. Accuracy is the evaluation metric.
  • QQP (Quora Question Pairs): Focuses on determining if two questions asked on Quora are semantically equivalent. The metrics for evaluation are accuracy and F1 score.
  • RTE (Recognizing Textual Entailment): Sentence entailment task. Accuracy is used as the metric.
  • STS (Semantic Textual Similarity Benchmark): Evaluates the degree of semantic similarity between two sentences. The primary metric is Pearson and Spearman correlation coefficients.
These tasks provide a broad evaluation of the model’s performance across various dimensions of NLU. The experiment results are shown in Table 2
@@
DeBERTa-base-h256 [13] 79.04 74.50 79.41/79.99 78.19 87.88 88.68 64.62 88.76 80.11
BERT+BPDec-base-h256 79.02 75.74 79.08/80.33 81.86 88.05 88.46 62.09 89.33 80.44
@@
DeBERTa-large [13] 87.23 85.62 87.01/87.54 87.75 93.14 91.17 76.53 94.84 87.87
BERT+BPDec-large 87.12 85.81 87.36/87.28 88.24 93.28 91.42 76.90 94.27 87.96
@@
In our finetuning experiments, each model was trained on the designated training set and subsequently evaluated on the development set. Table 2 illustrates that under identical conditions and with the same number of pretraining samples, BERT with BPDec enhancement consistently outperforms both BERT and DeBERTa models in terms of many tasks and the averaged score. Notably, MNLI, the largest dataset among these tasks, serves as a robust indicator of model performance. Our model competes closely with the similarly sized DeBERTa model, achieving slightly higher scores in the matched condition and comparable results in the mismatched one. However, it’s crucial to consider that the DeBERTa model’s encoder is approximately 17% larger in parameter size than BERT encoder, and it incurs a 32% higher cost in finetuning cost and a 29% increase in inference latency. Despite these disparities, BERT+BPDec matches or even surpasses DeBERTa in various benchmarks in base-h256, base and large level sizes. Moreover, BERT+BPDec achieves significant improvement over the original BERT at exactly the same finetuning and serving costs. The results tell us that this decoder only enhancement significantly boosts performance without incurring further additional costs. Building on this, we extended our testing to question answering tasks, further solidifying the conclusion that our work can yield substantial performance gains.
@@ 4.3 Performance on SQuAD Tasks
BERT+BPDec-base-h256 75.74 84.14 67.97 70.82
@@
DeBERTa-base [13] 83.75 90.78 77.87 80.94
BERT+BPDec-base 84.20 91.12 78.11 81.23
@@
BERT+BPDec-large 86.11 92.70 82.12 85.39
@@
The Stanford Question Answering Dataset (SQuAD) serves as a benchmark for evaluating question answering systems in NLP. It comprises two versions: SQuAD v1, focused on answerable questions from given passages, and SQuAD v2, which includes both answerable and unanswerable questions. Performance is measured using Exact Match (EM) and F1 Score, reflecting the precision and accuracy of the model’s responses.
The experiments conducted on both the v1 and v2 tasks, as presented in Table 3, followed a similar trend to those performed on the GLUE tasks: all models were trained on the training set and evaluated on the development set. While the results for BERT+BPDec on SQuAD v1 did not achieve the highest scores in the table, its performance on SQuAD v2, particularly for the base and base-h256 size model, was notably stronger than the original BERT. Especially, when considering that BERT+BPDec competes closely with, or even surpasses, the DeBERTa model, the improvement is significant. These results underscore BPDec’s efficiency and efficacy, showcasing its ability to deliver robust performance in question answering tasks, a key area in NLP, without additional overhead.
@@ 4.4 Ablation Study
The BPDec involves multiple modifications of architecture and randomization which introduce many parameters in the process. It is essential to clearly articulate whether each change is necessary and how much it contributes to the final result. Furthermore, understanding the influence of each major parameter on the outcome is also crucial. To further demonstrate the effectiveness of our method, we conducted several ablation experiments on selected GLUE tasks and SQuAD tasks with base-level sized models.
@@ 4.4.1 Number of MLM Decoder Layers
Our findings in Table 4 revealed that adding two decoder layers to the BERT-base model yielded the best results. Parallel experiments were conducted on the larger model as well, where adding four decoder layers to the BERT-large model proved most effective. These results led us to conclude that augmenting the model with additional decoder layers equivalent to one-sixth of the number of encoder layers seems to be the optimal strategy.
4.4.2 Decoder Layers with Attention Mask Relaxation
Decoder layer unmasking rates AX COLA MNLI-m/mm MRPC QNLI QQP SQuAD v2 EM/F1
@@ 4.4.2 Decoder Layers with Attention Mask Relaxation
Table 5: Comparative results on various tasks for a 12-layer BERT + 2-layer decoder model with different configurations regarding the application of attention mask restrictions during pretraining. The finetuning is applied to BERT encoder only. The configurations are denoted by lists, where each list specifies 2 random unmasking rates for the first and second decoder layers
In our ablation study shown in Table 5, we examined the effects of releasing attention mask restrictions across various configurations of the BPDec model during the MLM pretraining. Initially, we removed the attention mask entirely for the decoder, where we observed no positive impact on performance. However, a notable improvement emerged when the same restriction release was applied to a deeper layer of the BPDec decoder. Moreover, we found an optimal configuration when we gradually remove the mask, the first layer randomly unmask 50%, and the second decoder layer release the 50% attention mask left. This specific adjustment yielded the most significant performance enhancements. These results highlight the nuanced interplay between model architecture and attention mechanisms, emphasizing the potential of targeted modifications in MLM decoder to achieve better outcomes in complex language processing tasks. Additionally, we observed that the improvement on the SQuAD task after adding attention mask relaxation was significantly larger than the gains on other classification tasks. Considering that SQuAD tasks generally have longer and more complex contexts, this indirectly indicates that our improvement can enhance the model’s ability to understand context.
@@ 4.4.3 Effectiveness of Output Random Mix
In our ablation studies, we explored the integration of randomness into the model output before the Softmax layer with Masked Language Modeling Task. A particularly effective method was found is to randomly alternate between using outputs from the decoder and the encoder. We set our randomization strategy to only apply randomness along the (batch, seq) dimensions, maintaining the cohesiveness of individual output embeddings. This implies that an output embedding is sourced entirely from either the decoder or the encoder.
Furthermore, we experimented with varying mix ratios of decoder and encoder outputs. As shown in Table 6, the most effective ratio during pretraining was found to be 80% decoder output and 20% encoder output. This specific mix ratio led to the most balanced performance, underlining the importance of fine-tuning the balance between the two types of outputs to achieve optimal results in the finetuning phase. The 80% ratio for utilizing decoder output emerged as a reasonable choice. If this ratio were lower, it might result in insufficient training of the decoder, diminishing its potential benefits. Conversely, if the ratio were higher, it could undermine the intended effect of randomness.
@@ 4.5 Generalizability to Other BERT-like Models
To further demonstrate the effectiveness of BPDec, we applied the same methodology to ALBERT [17] and BERT-PreLN [16] models. For ALBERT, the BPDec does not share parameters with the ALBERT encoder, but the decoder layers within BPDec share parameters. For BERT-preLN, we followed the encoder’s structure and used pre-LayerNorm in BPDec. Experiments on the GLUE benchmark and SQuAD dataset showed consistent performance improvements across both models.
@@
BERT-PreLN+BPDec-base 85.42 85.08 85.55/85.86 88.75 91.71 90.77 78.04/80.92
Table 7: Comparative results of ALBERT and BERT-PreLN models with and without BPDec.
These additional experiments shown in 7 prove the effectiveness, adaptability, and generalizability of the BPDec method, highlighting its potential for broader application in various BERT-like architectures. The consistent improvements observed across different model types suggest that BPDec’s principles can be successfully transferred to enhance the performance of a wider range of models.
@@ 5 Conclusion
In this paper, we introduced BERT+BPDec, an enhanced version of the original BERT model, focusing on improvements with the Masked Language Modeling (MLM) Decoder used in pretraining. Key innovations include adding transformer block layers as MLM decoder, removing restrictions on attending to masked positions in the MLM decoder, and introducing a degree of output randomness by randomly mixing encoder and decoder outputs. Through rigorous evaluations on tasks such as MNLI, SQuAD, and RACE, BERT+BPDec model demonstrated superior performance over the original BERT and other state-of-the-art models, without any increasing computational complexity in further finetuning and inference. The ablation study confirmed the effectiveness of each modification, highlighting BPDec’s potential in advancing NLP efficiency and effectiveness. The optimal hyperparameter settings we provide for BPDec not only improve the model’s performance and efficiency but also contribute to sustainable computing by reducing energy consumption and emissions. This work has practical implications for real-world applications of language models and paves new avenues for future research on optimizing the masked language modeling pretraining process.

### v4

@@ 4.2 Performance on GLUE Tasks
  • AX (GLUE Diagnostic Task): Tests ability to understand linguistic phenomena such as logic, predicate-argument structure, and lexical semantics.
  • CoLA (Corpus of Linguistic Acceptability): Measures grammatical correctness.
  • MNLI-m/mm (Multi-Genre NLI - Matched/Mismatched): Assesses the ability to predict textual entailment across different genres.
  • MRPC (Microsoft Research Paraphrase Corpus): Focuses on identifying whether two sentences are paraphrases of each other.
  • QNLI (Question NLI): Involves determining whether a context sentence contains the answer to a question.
  • QQP (Quora Question Pairs): Focuses on determining if two questions asked on Quora are semantically equivalent.
  • RTE (Recognizing Textual Entailment): Sentence entailment task.
  • STS (Semantic Textual Similarity Benchmark): Evaluates the degree of semantic similarity between two sentences.
@@
DeBERTa-base-h256 [8] 79.04 74.50 79.41/79.99 78.19 87.88 88.68 62.62 88.76 79.89
BERT+BPDec-base-h256 79.06 75.74 79.39/80.33 81.86 88.05 88.96 62.59 89.33 80.59
@@
DeBERTa-large [8] 87.23 85.62 87.01/87.54 87.75 93.14 91.17 76.53 94.82 87.87
BERT+BPDec-large 87.22 86.01 87.76/87.49 88.24 93.28 91.42 76.90 94.85 88.16
@@
In our finetuning experiments, each model was trained on the designated training set and subsequently evaluated on the development set. Table 3 illustrates that under identical conditions and with the same number of pretraining samples, BERT with BPDec enhancement outperforms both BERT and DeBERTa models. Notably, MNLI, the largest dataset among these tasks, serves as a robust indicator of model performance. Our model competes closely with the similarly sized DeBERTa model, achieving slightly higher scores in the matched condition and comparable results in the mismatched one. However, it’s crucial to consider that the DeBERTa model’s encoder is approximately 17% larger in parameter size than BERT encoder, and it incurs 32% additional costs in finetuning and a 29% increase in inference latency. Despite these disparities, BERT+BPDec matches or even surpasses DeBERTa in various benchmarks in base-h256, base and large level sizes. Moreover, BERT+BPDec achieves significant improvement over the original BERT with consuming the exact same finetuning and serving costs. The results tell us that this decoder only enhancement significantly boosts performance without incurring additional costs. Building on this, we extended our testing to question answering tasks, further solidifying the conclusion that our work can yield substantial performance gains.
@@ 4.3 Performance on SQuAD Tasks
BERT+BPDec-base-h256 75.83 84.24 68.00 70.12
@@
DeBERTa-base [8] 83.75 90.78 77.87 81.04
BERT+BPDec-base 84.30 91.22 78.21 81.43
@@
BERT+BPDec-large 86.81 92.90 82.62 85.59
@@
The Stanford Question Answering Dataset (SQuAD) serves as a benchmark for evaluating question answering systems in NLP. It comprises two versions: SQuAD v1, focusing on answerable questions from given passages, and SQuAD v2, which includes both answerable and unanswerable questions. Performance is measured using Exact Match (EM) and F1 Score, reflecting the precision and accuracy of the model’s responses.
The experiments conducted on both the v1 and v2 tasks, as presented in Table 4, following a similar trend to those performed on the GLUE tasks: all models were trained on the training set and evaluated on the development set. While the results for BERT+BPDec on SQuAD v1 did not achieve the highest scores in the table, its performance on SQuAD v2, particularly for the base and base-h256 size model, is notably stronger than the original BERT and even surpases the DeBERTa model. These results underscore BPDec’s efficiency and efficacy, showcasing its ability to deliver robust performance in question answering tasks, a key area in NLP, without additional overhead.
@@ 4.4 Ablation Study
The BPDec involves multiple modifications of architecture and randomization which introduce many parameters in the process. It is essential to clearly articulate whether each change is necessary and how much it contributes to the final result. Furthermore, understanding the influence of each major parameter on the outcome is also crucial. To further demonstrate the effectiveness of our method, we conducted several ablation experiments on selected GLUE tasks and SQuAD tasks with base-level sized models. For these experiments, we strategically selected GLUE tasks that have larger datasets to ensure robustness and credibility in our results and to fit the data representations within a single table effectively. Tasks like RTE and SST, which have relatively smaller datasets were ommited. In contrast, tasks such as MNLI, with its substantial dataset of nearly 400,000 examples, were included. This selection criteria is designed to illustrate the impact of our modifications more clearly and designed to be fitted into a single table. By focusing on tasks with larger datasets, we can provide more significant insights, but can still validate the efficacy of our approach comprehensively.
@@ 4.4.1 Number of MLM Decoder Layers
Our findings in Table 5 revealed that adding two decoder layers to the BERT-base model yielded the best results. Parallel experiments were conducted on the larger model as well, where adding four decoder layers to the BERT-large model proved the most effective. These results led us to conclude that augmenting the model with additional decoder layers equivalent to approximately one-sixth of the number of encoder layers seems to be the optimal strategy.
4.4.2 Decoder Layers with Gradual Unmasking Attention (GUA)
GUA unmasking rates AX COLA MNLI-m/mm MRPC QNLI QQP SQuAD v2 EM/F1
@@ 4.4.2 Decoder Layers with Gradual Unmasking Attention (GUA)
Table 6: Comparative results on various tasks for a 12-layer BERT + 2-layer decoder model with different configurations on Gradual Unmasking Attention (GUA) in each layer regarding the unmasking rates during pretraining. The configurations are denoted by lists, where each list specifies 2 random unmasking rates for the GUA in the first and the second decoder layers
In our ablation study shown in Table 6, we examined the effects of releasing attention mask restrictions across various configurations of the BPDec model during the MLM pretraining. Initially, the complete removal of the attention mask for the decoder under the conventional setting showed no significant impact on performance. However, significant enhancements were observed when the Gradual Unmasking Attention (GUA) strategy was applied to a deeper layer of the BPDec decoder. The optimal performance was achieved with a specific implementation of GUA: in the first decoder layer, 50% of the positions were randomly unmasked, and in the second layer, the remaining 50% were fully revealed. This configuration of GUA yielded the most substantial improvements across our tests. These findings underscore the intricate relationship between model architecture and the dynamic application of attention mechanisms, highlighting the transformative potential of GUA in refining MLM decoders for complex language processing tasks. The Figure 4 is a heatmap illustrating the average attention weights over all attention heads in the last layer of a pretrained BERT+BPDec model. The visualization highlights how certain attention heads could attend to masked token embeddings, comparing the traditional masked attention mechanism with our proposed Gated Universal Attention (GUA) approach. Moreover, the enhancements on the SQuAD task, observed after integrating GUA, were significantly larger compared to those on other classification tasks. This indicates that GUA particularly boosts the model’s ability to process and understand longer and more complex textual contexts, affirming its efficacy in tasks requiring deep contextual comprehension.
Figure 4: Attention heatmap comparing masked attention and our GUA mechanism in the last layer of a pretrained BERT+BPDec model.
@@ 4.4.3 Effectiveness of Output Random Mix
In our ablation studies, we explored the integration of randomness into the model output before the Softmax layer with Masked Language Modeling Task. A particularly effective method we found is to randomly alternate between using outputs from the decoder and the encoder. We experimented with varying mix ratios of decoder and encoder outputs. As shown in Table 7, the most effective ratio during pretraining was found to be 80% decoder output and 20% encoder output. This specific mix ratio led to the most balanced performance, underlining the importance of fine-tuning the balance between the two types of outputs to achieve optimal results in the finetuning phase. The 80% ratio for utilizing decoder output emerged as a reasonable choice. If this ratio were lower, it might result in insufficient training of the decoder, diminishing its potential benefits. Conversely, if the ratio were higher, it could undermine the intended effect of randomness.
@@ 4.5 Generalizability to Other BERT-like Models
+BPDec 85.42 85.08 85.55/85.86 88.75 91.71 90.77 78.04/80.92
DeBERTa-base 85.43 82.74 85.51/85.86 88.48 92.17 91.18 78.21/81.43
+BPDec 85.43 83.31 85.49/85.91 88.56 91.98 91.21 78.60/81.56
Table 8: Comparative results of ALBERT, BERT-PreLN and DeBERTa models with and without BPDec. For DeBERTa-base, +BPDec means adding GUA and Random Mix to the original DeBERTa decoder.
To further demonstrate the effectiveness of BPDec, we implemented and applied the same methodology to ALBERT [13], BERT-PreLN [12] and DeBERTa [8] models. For ALBERT implementation, the BPDec does not share parameters with the ALBERT encoder, but the decoder layers within BPDec share parameters. For BERT-preLN, we followed the encoder’s structure and used pre-LayerNorm in BPDec, ensuring the structural parity between encoder and decoder. Experiments on the GLUE benchmark and SQuAD dataset shows consistent performance improvements across the models. For DeBERTa model, because the original model already has a similar decoder structure, we added GUA and applied random mix upon the original decoder.
These additional experiments shown in 8 prove the effectiveness, adaptability, and generalizability of the BPDec method, highlighting its potential for broader application in various BERT-like architectures. The consistent improvements observed across different model types suggest that BPDec’s principles can be successfully transferred to enhance the performance of a wider range of models.
@@ 5 Conclusion
In this paper, we introduced BERT+BPDec, an enhanced version of the original BERT model, focusing on improvements with the MLM Decoder used in pretraining. Key innovations include the additions of transformer blocks as an MLM decoder, the removal of restrictions on attending to masked positions via the novel Gradual Unmasking Attention (GUA), and the introduction of output randomness by randomly mixing encoder and decoder outputs. These enhancements allowed the BERT+BPDec model to demonstrate superior performance over both the original BERT and other state-of-the-art models in rigorous evaluations on tasks such as MNLI, SQuAD, and RACE, without increasing computational complexity during further finetuning and inference. The BPDec not only facilitates improved performance and efficiency but also contributes to sustainable computing by reducing energy consumption and emissions. Moreover, the ablation study underscored the effectiveness of each modification and the optimal hyperparameter settings are provided. This work has practical implications for real-world applications of language models and paves new avenues for future research on optimizing the language modeling. While our research currently only focuses on enhancing BERT and similar MLM-based models, we are enthusiastic about its potential to inspire adaptations in general language modeling that could benefit models like GPT and T5. Our work underscores a commitment to reducing computational waste and emissions across models and their applications, setting a precedent for environmentally conscious innovations in natural language processing.